Ticket Number: HEA-NOC/20071130-4 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20071130 10:03 UTC Problem Start: 20071203 08:00 UTC
Ticket Closed: 20080107 13:55 UTC Problem End: 20071203 10:00 UTC
It is planned to migrate our INEX Lan #1 link to the new IP equipment in
Citywest on Monday 3rd December.
This work is contingent on the successful completion of the MNS services
migration, as per ticket HEA-NOC/20071126-7
Traffic to INEX peers will reroute via Lan #2 (where possible) or the
general internet for the duration of the physical repatching and
reconfiguration. There may be a brief interruption to service while
routing protocols reconverge.
20071130 09:59 DW Open ticket.
20071201 13:40 DW MNS work successful, so beginning preparation work for
this change. This preparation work should be non-service affecting.
20071201 15:56 DW Preconfiguration complete. Maintenance to proceed at
08:00 on Monday.
20071203 12:30 DW Maintenance took place. BGP sessions were shut down
beginning 08:21, and brought up on the new router between 08:49-09:23.
Traffic is flowing as expected and we continue to monitor.
20080107 13:54 DW Closing ticket. There will be a swapout of equipment in
the coming weeks that will be tracked separately.
Time to Fix:
Scheduled 2 hours.
This ticket can be monitored at http://www.hea.net/tickets/20071130-4
HEAnet Limited. Registered in Ireland, No. 275301.